I need math help :/ Stacey's student loan totaled $12000. One of her loans was at 6.5% simple interest and the other at 7.2%. After one year, Stacey owed 811.50 in interest. What was the amount of each loan?I don't understand!
Answer by mananth(16946)

Part I 6.50% per annum ------------- Amount invested =x
Part II 7.20% per annum ------------ Amount invested = y
12000
Interest----- 811.50 for one year

Total investment
x + y= 12000 -------------1
Interest on both investments
6.50% x + 7.20% y= 811.5
Multiply by 100
6.5 x + 7.2 y= 81150.00 --------2
Multiply (1) by -6.5
we get
-6.5 x -6.5 y= -78000.00
Add this to (2)
0 x 0.7 y= 3150
divide by 0.7
y = 4500
Part I 6.50% $ 7500
Part II 7.20% $ 4500

CHECK
7500 --------- 6.50% ------- 487.50
4500 ------------- 7.20% ------- 324.00
Total -------------------- 811.50

m.ananth@hotmail.ca